OpenStudy (anonymous):

The polynomial given below has__ roots(s). 3x^2+4x+3 A. one negative and one positive B. two negative C. two complex D. two positive

OpenStudy (mathstudent55):

Do you know the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Do you know the discriminant of the quadratic formula? It's the radicand of the quadratic formula.

OpenStudy (anonymous):

nope

OpenStudy (mathstudent55):

The discriminant is the part: \(b^2 - 4ac\) Plug in the numbers into the discriminant. Then if the discriminant: 1. = 0, there is one real solution. 2. < 0, there are two complex solutions 3. > 0, there are two different real solutions

OpenStudy (anonymous):

i got 20

OpenStudy (mathstudent55):

\(3x^2+4x+3= 0\) compare with \(ax^2 + bx + c = 0 \) a = 3, b = 4, c = 3 \(b^2 - 4ac = 4^2 - 4(3)(3) = 16 - 36 = -20 \) The discriminant is -20, which is negative, or less than zero. What can you conclude about the roots of the polynomial?

OpenStudy (anonymous):

B

OpenStudy (mathstudent55):

Read again what I wrote above. I copied it below to make it easier for you to find it. Your discriminant is -20, a negative number. Then if the discriminant: 1. = 0, there is one real solution. \(\color{red}{\bf{2. < 0, there~ are~ two~ complex~ solutions}}\) 3. > 0, there are two different real solutions

OpenStudy (anonymous):

thanks bratha

OpenStudy (mathstudent55):

Answer is C. Wlcm.